• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 09:43:50 on April 06, 2008

    Entry number 230330

    This entry is a followup to: 230321

    keyword=re: helicity

    The problem with helicity is probably the same as what affected scalers yesterday (see my
    halog 230053). If that's right, then fixing it in the datastream is simply a matter of reversing
    "0" and "1" for the gate (MPS) in software. This gate tells when the helicity is valid. Invalid
    helicity is 0, normal helicity is +/-1. The helicity is invalid during the 500 usec transition time
    when it is changing, so the peak at zero should be about 1.5% of data, but now it seems
    flipped.

    After I fixed the scalers yesterday I found out that nobody had deliberately flipped MPS, and
    that at 3 pm on Fri Apr 4 Bryan had recabled MPS to avoid a bad output. I think we should
    check this again to be sure we really have MPS and that we understand what is going on. I'll
    come in a little later to check.

    A copy of this log entry has been emailed to: moffit@jlab.org,vasulk@jlab.org